can you help me with this question :
A merchant has some coffee worth 80 cents and some worth 90 cents per pound. How many pounds of each should he use to make 100 pounds of a mixture worth 88 cents per pound?

Answer by mananth(16946)

Price ---------------- quantity
coffee type I 0.90 ---------------- x pounds
Coffee type II 0.80 ------ 100 - x pounds
Mixture 0.88 ---------------- 100
100
0.9 x + 0.8 ( 100 - x ) = 88
0.9 x + 80 - 0.8 x = 88
0.9 x - 0.8 x = 88 - -80
0.1 x = 8
/ 0.1
x = 80 pounds 90 cents coffee type I
20 pounds 80 cents Coffee type II

m.ananth@hotmail.ca